A 127 mL bubble of hot gases at 213 ∘C and 1.78 atm is emitted from an active volcano.

Part A

What is the final temperature, in degrees Celsius, of the gas in the bubble outside the volcano if the final volume of the bubble is 136 mL and the pressure is 0.770 atm , if the amount of gas does not change?

Express your answer using two significant figures.
Answer #1

Given:

Pi = 1.78 atm

Pf = 0.770 atm

Vi = 127 mL

Vf = 136 mL

Ti = 213.0 oC

= (213.0+273) K

= 486 K

use:

(Pi*Vi)/(Ti) = (Pf*Vf)/(Tf)

(1.78 atm*127 mL)/(486.0 K) = (0.77 atm*136 mL)/(Tf)

Tf = 225 K

= (225 - 273) oC

= -48 oC

Answer: -48